OpenStudy (anonymous):

Write the result in standard form a+bi (-1+4i)(1+i)

OpenStudy (radar):

-1-i+4i+4i^2 -5+3i

OpenStudy (unklerhaukus):

\[(-1+4i)(1+i)\] \[=-1-i+4i+4i^2\] \[=-1+3i-4\] \[=-5+3i\] \[a=-5,~~ b=3\]